Mathematics 244 – Linear Algebra Definitions Quiz 3 – February 9, 2007

Your Name:

A) (5) Define: The mapping T : R^{n} → R^{m} is linear.

B) (5) True or false: If an n × n matrix A is invertible, then the system of linear equations Ax = b has a unique solution for every b ∈ R^{n}. Explain (briefly!)